Calculate the magnitude of the orbital magnetic field experienced by a 2p-electron in hydrogen
1
Expert's answer
2020-12-15T11:46:43-0500

The magnitude of the orbital magnetic field experienced by a 2p-electron in hydrogen is


"B=\frac{\mu_0Ze\vec{L}}{4\pi r^2}=0.3\text{ T}."